# An X-ray absorption spectrum database for iron-containing proteins

### v1

@@ Abstract
Earth-abundant iron is an essential metal in regulating the structure and function of proteins. This study presents the development of a comprehensive X-ray Absorption Spectroscopy (XAS) database focused on iron-containing proteins, addressing a critical gap in available high-quality annotated spectral data for iron-containing proteins. The database integrates detailed XAS spectra with their corresponding local structural data of proteins and enables a direct comparison of spectral features and structural motifs. Utilizing a combination of manual curation and semi-automated data extraction techniques, we developed a comprehensive dataset from an extensive review of literature, ensuring the quality and accuracy of data, which contains 437 protein structures and 1954 XAS spectrums. Our methods included careful documentation and validation processes to ensure accuracy and reproducibility. This dataset not only centralizes information on iron-containing proteins but also supports advanced data-driven discoveries, such as machine learning, to predict and analyze protein structure and functions. This work underscores the potential of integrating detailed spectroscopic data with structural biology to advance the field of biological chemistry and catalysis.
Background & Summary
X-ray Absorption Spectroscopy (XAS) is a synchrotron-based technique that provides insights into the local chemical environment of an element. In terms of metalloproteins, XAS can be used to study the local environment of metal ions which are often present in the active site of proteins, and are essential for understanding their structure and function[1, 2, 3]. Understanding the molecular structure from XAS can accelerate the development of catalysts for numerous industrial and environmental applications. However, current methods for XAS data analysis are time-intensive and laborious. Recently, data-driven deep learning methods have demonstrated great potential in materials science for predicting the properties of new materials, optimizing reaction conditions, and designing novel catalysts[4, 5, 6, 7, 8]. These methods can analyze vast amounts of data quickly and identify patterns that might not be apparent through traditional methods such as predicting the activity and stability of catalysts for hydrogen production[9].
The integration of deep learning with XAS data can provide deeper insights into the local chemical environments of catalysts, enabling the prediction of catalytic performance and guiding the design of more efficient catalysts[10]. However, the effectiveness of these data-driven methods depends on the availability of comprehensive and high-quality datasets. This highlights the need for a comprehensive protein structure-XAS database. Such a database would supply the necessary data for training and evaluating deep learning models, enabling the rapid and automated analysis of protein structures and their catalytic properties.
Substantial databases, such as XASLIB[11] the Materials Project[12], offer extensive spectral data for inorganic compounds. These resources are important for advancing material sciences and are widely used for predicting material properties, designing new materials, and understanding fundamental processes[13, 14, 15, 16]. Despite these advances, there remains a significant gap in protein-focused XAS data. This study fills this gap by creating a reliable dataset of iron-containing protein structures and their corresponding XAS spectra, manually curated from the literature.
Here, we present the first comprehensive iron-containing protein structure-XAS database, addressing a critical need in the field. Our focus on iron is motivated by its essential role in numerous biological processes and the complexity of its coordination chemistry. Including other elements would dilute the specificity and introduce variability that could complicate initial analyses. Unlike existing databases like the Protein Data Bank (PDB)[17, 18] and the Cambridge Structural Database (CSD or CCDC)[19], which primarily focus on three-dimensional structures of proteins and small molecules respectively, our database integrates detailed XAS spectra with protein structural data, enabling direct comparison of XAS features with structural motifs.
In our work, we gathered Fe K-edge XAS data, together with their reported three-dimensional local iron environment of corresponding protein structures from the literature, resulting in a total of 437 protein structures and 1652 XAS spectrums obtained from 573 research papers published between 2007 and 2023. The dataset includes XANES and EXAFS spectra associated with various iron-containing proteins, spanning a wide range of experimental conditions and measurement techniques. This comprehensive dataset provides a comprehensive foundation for subsequent data-driven analysis and modeling efforts.
Methods
@@
The schematic overview of the data extraction pipeline is shown in Figure 1. Firstly, we constructed searching keywords to identify relevant literature. The search was conducted using combinations of keywords from two sets: {XAS, XANES, XAFS, EXAFS} and {Metalloprotein, Protein, Enzyme, Fe-iron}. Each search term consisted of a pair of words, one from each set, such as "XAS Protein" or "EXAFS Metalloprotein". These keywords were generated by combining terms from two distinct keyword sets, as illustrated in the figure. The searching keywords constructed from these two sets were used to retrieve the literature from various publishers, including Springer Nature, AAAS Science, American Chemical Society, Elsevier, Royal Society of Chemistry, Electrochemical Society, Wiley, MDPI, and others. To locate relevant literature, we utilized either the publishers’ search APIs or conducted manual searches under copyright permission. We excluded duplicate results, ending up with 20,915 articles (Table 1). After that, we manually selected papers that include both protein structure and the absorption spectra of iron elements, with the simple selection criterion of mentioning the protein and including at least one Fe absorption spectrum. Eventually, we obtained 573 articles that met these criteria, which were downloaded in PDF format to allow human experts to extract the necessary dataset plots. After gathering the literature, the human experts annotated the text and XAS plot information, converting it into digitized data samples. The extracted dataset was then refined by removing the low-quality samples, poorly presented XAS and structures, and outdated annotations. Finally, the extracted dataset was constructed through a combination of automatic searching and manual data extraction and cleaning. Each data sample in our dataset consists of three main parts: (1) The entire or local protein structure (The first-coordination sphere of the element of interest), (2) The protein’s corresponding Fe K-edge XAS spectra, and (3) the basic information of papers from which the protein structure and XAS spectrum were derived.
@@
Extracting XAS data from the retrieved literature is a crucial step in constructing our database and further contributing to the analysis of protein structure and function. XAS is divided into two regions: The near-edge spectra also known as X-ray absorption near-edge structure (XANES) and the extended X-ray absorption fine structure (EXAFS), each of them providing specific information on the element studied. To extract numerical data points from the published XAS plots, we utilized the open-source data annotator WebPlotDigitizer V4[20]. To guarantee the precision and reliability of data extraction we followed the steps below:
In the process of extracting X-ray Absorption Spectroscopy (XAS) data from published plots, we employ a meticulous digitization approach using the WebPlotDigitizer tool. Initially, we manually mask the spectral lines using the pencil tool, ensuring precise alignment with the plot data. The color recognition distance is set variably at 2, 5, or 10 pixels based on the clarity and overlap of the plotted lines; clearer plots are annotated at a denser resolution of 1 point per 2 pixels, while more complex, fuzzy plots require a broader setting of 1 point per 10 pixels. This flexibility in setting allows us to capture data with high fidelity, respecting the original plot’s integrity without interpolation during the initial annotation phase.
Post-annotation, to standardize the spectral data for consistent analysis and comparison, we interpolate all spectra to a uniform length of 100 data points, based on the mean distribution of spectra lengths observed across the dataset. As Figure 2 shown, the spectra lengths of XANES and EXAFS range from 0 to 600, and their peak values are located at around 100. Although XANES length distribution is more concentrated than EXAFS, it still indicates considerable variability in spectrum length.
In particular, we also preserve our dataset’s original extracted data points. This dual provision enables researchers to select between the original or interpolated data depending on the requirements of their specific analytical applications, thus offering flexibility while minimizing potential artifacts introduced by changes in data resolution.
What’s more, in the application of XAS in physical analysis or machine learning, the calibration energies critical for the accuracy of XAS measurements. In our dataset, the calibration energies are manually extracted from sentences within each annotated paper. However, it’s important to note that not all sources uniformly report this value. Specifically, some papers mention the use of iron foil as a standard without detailing the exact calibration energy, while others omit this information entirely. In our efforts to maintain transparency and provide a comprehensive dataset, we have documented each instance of missing or incomplete calibration information. Out of the dataset, 227 entries lack explicitly reported calibration energies. These cases are clearly labeled in the documentation files as shown in Fig 5(b), enabling researchers to easily identify and account for potential inconsistencies or gaps in data calibration when using the dataset for their scientific investigations.
@@
A comprehensive analysis of protein structure information was conducted through meticulous manual review of each paper from the database. The focus was specifically on iron-containing proteins and small iron-containing molecules of interest, which were characterized using iron (Fe) K-edge X-ray absorption spectroscopy. While other sample types, including tissues, blood, soil, water, and plant imaging, were noted in the comments section, the primary emphasis remained on iron-proteins and structure-solved small molecules. The extracted information was categorized into two distinct groups: near-edge spectra and extended X-ray absorption fine structure (EXAFS). Within these groups, samples were annotated with corresponding standard references. Additionally, Protein Data Bank (PDB) and Cambridge Crystallographic Data Centre (CCDC) codes were extracted from papers reporting structural data for proteins and small molecules, respectively.
Data Records
The dataset developed in this study is hosted on https://airscker.github.io/XDIP. The format of the data records is detailed in Figure 3, which illustrates that each data sample consists of three parts: literature metadata, Fe K-edge XAS data, and the protein’s local structure containing Fe.
The literature metadata in each data record includes the Digital Object Identifier (DOI), the title of the research paper, and the XAS absorbing element, which is Fe in this work. These metadata serve as a quick reference for identifying the original research papers and their primary findings. Since most papers cover multiple sets of iron K-edge XAS and corresponding structures, we have stacked multiple absorption spectra and protein structures to capture all relevant details in the literature. Additionally, in some cases, the spectrum may cover only XANES or EXAFS, resulting in records that contain only one type of spectrum.
The protein structure includes atom types, their spacial coordinates, and their relationships. We extracted the protein’s structure or Fe-element neighborhood local structure through two methods: manual annotation of structure information and accessing data from public protein structure databases. The manually extracted protein (local) structures were derived through the following steps:
In contrast to the manually extracted protein local structures, the extraction of local structures from open databases is more straightforward. In these cases, we simply include the URL for downloading the structure or the local path to access the saved structure files. Alternatively, the structure can be directly derived from the protein’s SMILES representation. Table 2 provides a detailed description of the data format, including the description of every key label, whether it is mandatory as well as the data type of the key labels.
In the final compilation of our dataset, we assembled a total of 437 protein local structures and 1652 corresponding XAS spectra, consisting of 1283 XANES and 369 EXAFS spectrums. The sample size was determined based on the availability and relevance of data from the literature, ensuring a comprehensive representation of diverse experimental conditions. Certain papers were excluded during the data curation process due to insufficient quality. Specifically, we omitted papers where figures lacked proper axis labels or annotations, or where the spectral curves were too indistinct to be reliably analyzed. Additionally, documents featuring protein samples not pertinent to our research, such as tissue samples, were excluded to maintain the relevance and specificity of the dataset to iron-containing proteins. These exclusions were essential to preserve the integrity and quality of the final dataset.
Technical Validation
@@
To ensure the accuracy of our annotated XAS spectra, we conducted a comparison between the annotated data and the raw data. Each spectrum was annotated by four experts independently, and the results were evaluated using Mean Squared Error (MSE) to measure the deviation of annotated values from the raw data. The steps involved in conducting the comparison are outlined below:
  • Data Source and Generation. Since most scientific papers do not provide raw data for corresponding spectrum plots, the validation data were randomly chosen from the Materials Project[12], focusing on the Fe element. We generated 50 spectrum plots for XANES and 25 for EXAFS. These plots had various shape ratios: 10:8, 10:6, 6:6, and 8:6, with the number of curves in a single plot ranging from 1 to 4. The line styles included dots, dashes, and solid lines, designed to simulate actual plots found in scientific papers.
  • Anonymous Annotation Workload Insertion. To ensure the accuracy of annotation and eliminate the possible bias of extracted spectrum points, we inserted these generated plots into every expert’s XAS annotation workload anonymously.
  • Data Alignment. Before calculating the MSE between the annotated and raw data, it was necessary to align the x-values of the spectrum data points. Due to the variability among experts, the annotated data may have different lengths and positions. To maintain consistency, the annotated spectra were linearly interpolated to have the same x-values as the raw data.
  • MSE Calculation: The MSE values were calculated for each spectrum and averaged across the entire dataset. The low average MSE indicates a high level of accuracy in the annotations when compared to the raw data.
Table 3 displays the MSE values for XANES and EXAFS spectra annotations by four experts, reflecting the challenges of annotating complex spectral data. The XANES MSE values are relatively low, ranging from 0.0007 to 0.0521, which suggests high accuracy in annotations given the simpler nature and lower fluctuation of XANES spectra within the raw data range of 0 to 1.6. In contrast, the EXAFS MSE values are considerably higher, from 0.0928 to 0.1015, highlighting the difficulty in accurately annotating these spectra due to their greater complexity and variability. The average MSE for XANES at 0.0176 and for EXAFS at 0.0958 further underscores the impact of spectral complexity on annotation accuracy, with EXAFS presenting a more challenging task for precise annotations compared to XANES. This distinction between the two types of spectra demonstrates how the intricacies of spectral data significantly influence the reliability and consistency of expert annotations in scientific research involving X-ray absorption spectroscopy.
Figure 4 illustrates the distribution of MSE values for three experts annotating XANES and EXAFS spectra for every sample. In Figure 4(a), values are generally low and range from $10^{-5}$ to $10^{-3}$, indicating high accuracy and the relative simplicity of XANES spectra. Conversely, Figure 4(b) shows a spread of larger values tightly clustered around $10^{-1}$, reflecting the complexity and variability inherent in EXAFS data. This larger value underscores the challenges posed by the more fluctuating patterns of EXAFS spectra compared to the more stable XANES spectra, resulting in variable annotation accuracy.

### v2

# XDIP: A Curated X-ray Absorption Spectrum Dataset for Iron-Containing Proteins

@@ Abstract
Earth-abundant iron is an essential metal in regulating the structure and function of proteins. This study presents the development of a comprehensive X-ray Absorption Spectroscopy (XAS) database focused on iron-containing proteins, addressing a critical gap in available high-quality annotated spectral data for iron-containing proteins. The database integrates detailed XAS spectra with their corresponding local structural data of proteins and enables direct comparison between spectral features and structural motifs. Utilizing a combination of manual curation and semi-automated data extraction techniques, we developed a comprehensive dataset via extensive literature review, ensuring the quality and accuracy of data, which contains 437 protein structures and 1954 XAS spectrums. Our methods included careful documentation and validation processes to ensure accuracy and reproducibility. This dataset not only centralizes information on iron-containing proteins but also supports advanced data-driven discoveries, such as machine learning, to predict and analyze protein structure and functions. This work underscores the potential of integrating detailed spectroscopic data with structural biology to advance the field of biological chemistry and catalysis.
1 Background & Summary
X-ray Absorption Spectroscopy (XAS) is a synchrotron-based technique that reveals the local chemical environment of specific elements. For metalloproteins, XAS is particularly powerful for probing the coordination and electronic structure of metal ions that are often located at protein active sites and play key roles in catalysis and functiontangcharoen2014synchrotron ; rawat2020structural ; tangcharoen2019synchrotron . Interpreting XAS data can thus accelerate our understanding of protein mechanisms and the development of catalytic systems for industrial and environmental applications. However, the analysis of XAS data remains time-consuming and highly manual.
In recent years, data-driven approaches—especially deep learning—have shown significant promise in materials science, enabling property prediction, reaction optimization, and the design of novel compoundsbutler2018machine ; schmidt2019recent ; torrisi2020random ; guda2021understanding ; kartashov2021xas . These models can extract meaningful patterns from large datasets, achieving impressive results in tasks like predicting catalyst stability and activityfeng2021egemm . Integrating such methods with XAS can deepen our understanding of element-specific environments and accelerate materials discovery.
The effectiveness of deep learning, however, critically depends on the availability of large, high-quality datasets. While substantial XAS repositories like XASLIBxaslib and the Materials Projectjain2013commentary provide extensive data for inorganic compounds, they lack detailed spectroscopic data for biologically relevant molecules. These resources have greatly advanced materials informaticszhou2004first ; wang2007first ; waroquiers2020chemenv ; horton2019high , yet a significant gap persists in the realm of protein-focused XAS.
To address this, we introduce the first curated database of iron-containing protein structures paired with their corresponding XAS spectra. Iron was selected as the focal element due to its abundance in nature, central role in biological catalysis, and the complexity of its coordination chemistry. Including other metals at this stage would compromise specificity and introduce variability, limiting the ability to draw meaningful structure-spectrum correlations.
Unlike existing repositories such as the Protein Data Bank (PDB)berman2000protein ; zardecki2022pdb and the Cambridge Structural Database (CSD/CCDC)groom2016cambridge , which focus on structural information alone, our database couples Fe K-edge XAS spectra with detailed local structural annotations. This integration enables direct alignment between spectral features and structural motifs, supporting both mechanistic insights and machine-learning-driven discovery.
Our final dataset comprises 437 iron-containing protein structures and 1652 associated XAS spectra (including 1283 XANES and 369 EXAFS), all manually curated from 573 peer-reviewed articles published between 2007 and 2024. The samples span a diverse set of proteins, experimental conditions, and measurement protocols. This curated collection establishes a critical foundation for automated, high-throughput structure–spectrum modeling in protein chemistry and bioinorganic catalysis.
2 Methods
@@
We constructed the dataset using a semi-automated pipeline designed to maximize both coverage and quality (Figure 1). We began with a comprehensive literature search across major scientific publishers to identify articles reporting both iron-containing protein structures and corresponding Fe K-edge X-ray Absorption Spectroscopy (XAS) data. This process yielded 573 relevant publications for manual analysis.
Each article underwent a rigorous two-stage curation process by human experts: (1) digitizing XAS spectra from published figures, and (2) annotating the associated local protein structures and metadata based on textual descriptions. We then refined the dataset by removing low-quality entries or samples lacking sufficient documentation. Each finalized data sample in our database comprises three core components: (1) the local atomic structure surrounding the iron center, (2) the Fe K-edge XAS spectrum, and (3) metadata from the original publication. Full details of the search keywords, inclusion criteria, and curation protocol are provided in Supplementary Section A.1.
Figure 1: Schematic overview of the dataset construction pipeline. (a) Literature retrieval using keyword combinations from two sets (shown at the top) to search across multiple scientific databases. (b) Human expert processing workflow: relevant protein structures and spectra are identified and digitized, then curated into structured formats. (c) Final data sample: each entry includes the local protein structure, Fe absorption spectrum, and metadata from the source paper.
@@
We extracted numerical spectral data from figures using the open-source tool WebPlotDigitizer. Expert annotators manually digitized two key spectral regions: X-ray Absorption Near-Edge Structure (XANES) and Extended X-ray Absorption Fine Structure (EXAFS). To ensure high fidelity with the original plots, we carefully calibrated the plot axes and manually traced the spectral curves.
To enable compatibility with machine learning workflows, all spectra were interpolated to a uniform length of 100 data points, a choice based on the average distribution of spectrum lengths across the dataset. For transparency and flexibility, both the original digitized data and the interpolated 100-point versions are included. Further details on digitization settings and metadata documentation—including treatment of missing calibration energies—are provided in Supplementary Section A.2.
@@
Protein structural data were manually annotated from each paper, with a focus on iron-containing proteins and small iron-containing molecules characterized using Fe K-edge XAS. While other sample types—such as tissues, blood, soil, and plant imaging—were sometimes encountered, these were noted in the comments section and excluded from primary analysis.
We categorized the extracted samples based on spectral type (XANES or EXAFS) and annotated them with available standard references. Structural information included Protein Data Bank (PDB) or Cambridge Crystallographic Data Centre (CCDC) identifiers when available. When structural information was not publicly accessible, we manually extracted it from figures and textual descriptions, focusing on the local atomic environment around the iron center.
3 Data Records
Our final dataset consists of 437 unique iron-centered local protein structures and 1652 associated XAS spectra, including 1283 XANES and 369 EXAFS records. The complete dataset is publicly available at https://airscker.github.io/XDIP.
Each data record is a self-contained unit suitable for machine learning applications and includes three core components (illustrated in Figure 2): (1) literature metadata, (2) Fe K-edge XAS spectrum, and (3) the local atomic structure surrounding the iron center. The metadata component includes the source paper’s title and DOI to ensure full traceability. Since a single publication may contain multiple experiments, each record captures all distinct spectra and structures reported within that paper.
The local structure is formatted to support graph-based machine learning, including a list of atoms, their 3D Cartesian coordinates, and an adjacency matrix representing chemical bonds. This representation is further enriched with bond lengths and, where available, bond angles to offer a more complete geometric description. Structural data were sourced in two ways: (1) directly from public repositories such as the Protein Data Bank (PDB) or Cambridge Structural Database (CCDC), or (2) manually reconstructed from figures and descriptions in the literature. In some cases, structures were also retrieved or generated from SMILES representations.
To ensure data integrity and relevance, we applied strict curation criteria throughout the pipeline. Papers were excluded if figures lacked sufficient annotation or if the spectra were too noisy or incomplete for reliable digitization. Moreover, we maintained a focused scope on iron-containing proteins by excluding samples from other material classes (e.g., tissues, soil, or plant matter). A detailed description of the data schema—including field labels, types, and requirements—is provided in Supplementary Section A.3.
Figure 2: Overview of a structured data sample. Multiple records can be extracted from a single publication. (a) Metadata, including the paper’s DOI, title, and absorbing element (Fe in this dataset). (b) Spectrum data, divided into near-edge (XANES) and extended (EXAFS) regions. (c) Local structure, either retrieved from public databases (e.g., SMILESweininger1988smiles ; weininger1989smiles ; weininger1990smiles ) or manually annotated. Structural information includes an adjacency matrix, atom list, Cartesian coordinates, bond lengths, bond angles, and optional notes.
4 Technical Validation
@@
To assess the accuracy of the annotated XAS spectra, we performed a quantitative comparison between expert annotations and reference (ground-truth) spectra. Each spectrum was independently annotated by four experts, and the deviation from the reference data was measured using Mean Squared Error (MSE). The validation protocol included the following steps:
  • Data Source and Generation. Since most publications do not provide raw data for their spectrum plots, we generated synthetic validation spectra using Fe-based XANES and EXAFS data from the Materials Projectjain2013commentary . We produced 50 XANES and 25 EXAFS plots with varied shape ratios (10:8, 10:6, 6:6, 8:6), curve counts (1–4), and line styles (dots, dashes, solid), simulating real-world publication styles.
  • Anonymous Annotation Insertion. These synthetic plots were anonymously inserted into each expert’s annotation workload to avoid bias and to assess natural annotation performance.
  • Data Alignment. Because manual digitization may result in inconsistent data lengths and misaligned x-axes, we linearly interpolated all annotated spectra to match the x-values of the reference spectra.
  • MSE Calculation. We computed MSE for each spectrum and averaged results across experts. Low average MSE values indicate high annotation fidelity.
Table 1: Mean Squared Error (MSE) values of each expert for XANES and EXAFS.
Expert 1 Expert 2 Expert 3 Expert 4 Average
XANES MSE 0.0161 0.0014 0.0007 0.0521 0.0176
EXAFS MSE 0.0946 0.0928 0.1015 0.0944 0.0958
(a) XANES MSE distribution
(b) EXAFS MSE distribution
Figure 3: Distribution of annotation MSE values across experts for XANES (a) and EXAFS (b) spectra.
Table 1 shows that annotations of XANES spectra achieved lower MSEs (0.0007–0.0521), reflecting their smoother shape and simpler features. In contrast, EXAFS annotations showed higher MSEs (0.0928–0.1015), attributable to their more complex and oscillatory structure. This trend is reinforced by the histograms in Figure 3, where XANES errors are clustered between $10^{-5}$ and $10^{-3}$, while EXAFS errors concentrate around $10^{-1}$. These results highlight the increased difficulty and reduced consistency of EXAFS annotation due to its higher frequency content and structural sensitivity.